# Bifurcations and canards in two coupled FitzHugh–Nagumo equations July 25, 2025

### v1

@@ 5. Canards and mixed-mode oscillations
which is not zero and less than 1/4 for an open set of parameters $(k,b,c)$, since $\varphi^{\prime\prime}(x_{2}^{*})=-6x_{2}^{*}\neq 0$. ∎
Figure 6. Canard transient on a solution of (2) near a folded node, parameters $b=0$, $c=-0.519935054$, $k=1$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(-1.5,2,{\varphi(-1.5),\varphi(2)}\right)$. (A) - Time course for $y_{1}(t)$ (blue) and $y_{2}(t)$ (orange). (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) Projection of the trajectory (red) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the red dot, equilibria on the white dots. (D) - Projection of the trajectory (purple) on the $\left(x_{2},y_{1},y_{2}\right)$ space, initial condition on the blue dot, purple critical manifold $C_{0}$. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, conventions as in (B). (F) - Time course for $x_{1}(t)$ (blue) and $x_{2}(t)$ (orange).
The time rescaling we have used reverses time orientation when $\left(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*})\right)$ is in the attracting part of $C_{0}$ where $\varphi^{\prime}(x_{2}^{*})<0$, so when $(x_{1}^{*},x_{2}^{*})$ is either a saddle or an unstable node for (7) there is at least one trajectory of (2) that goes across $\left(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*})\right)$ generating a canard. This is a consequence of Lemma 2.3 of [25], where it is shown that canards correspond to, at least, one negative eigenvalue of (7). See also Guckenheimer & Haiduc [12] and Guckenheimer [13]. Examples are shown in Figures 6, 7 and 8.
In the first of these examples, before the trajectory approaches a synchronous periodic orbit it makes some transient small oscillations around the fold point that are visible in Figure 6 (E). In this example there is a folded node, see Table 1.
Note that from equation (10) it follows that for any given $k\neq 0$ and $b\in{\mathbf{R}}$ there is a value of the parameter $c\in{\mathbf{R}}$ for which $\det DH(x_{1}^{*},x_{2}^{*})=0$. This corresponds to saddle-node bifurcations. According to [17] this gives rise to mixed-mode oscillations like those shown in Figures 7 and 8, i.e., trajectories that combine small oscillations and large oscillations of relaxation type, both recurring in an alternating manner, cf. [17].
@@ 5.1. The case $b=0$
More information on the equilibria of the equation (7) may be obtained if we make the simplifying assumption $b=0$. Indeed the expression of (1) in the particular case $b=0$ has been used in the analysis of two FHN coupled in the fast equations by Pedersen et al. [22] and with double coupling fast–to–fast and slow–to–slow by Krupa et al. 2014 [17]. Thus this particular case is interesting for comparing the outcomes of different types of couplings.
@@
$A_{1}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2}):\ x_{1}>2/\sqrt{3},\quad x_{2}>2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}$
$A_{2}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2}):\ x_{1}<-2/\sqrt{3},\quad x_{2}>2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}$
$A_{3}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2}):\ x_{1}<-2/\sqrt{3},\quad x_{2}<-2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}$
$A_{4}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2}):\ x_{1}>2/\sqrt{3},\quad x_{2}<-2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}.$
@@
In Table 1, we present sufficient conditions on $c$ for the existence and classification of a folded equilibrium of (7) with $b=0$ on the components with $x_{2}^{*}=\pm 2/\sqrt{3}$ of $\partial A_{i}^{*}$, $i=1,\ldots,4$, the boundary of the attracting region of the critical manifold. The analysis for folded equilibria with $x_{2}^{*}=\pm 2/\sqrt{3}$ follows by symmetry. We present the computations for the cases where the folded equilibrium $\left(x_{1}^{*},x_{2}^{*}\right)$ lies on the boundaries of $A_{1}^{*}$ and $A_{2}^{*}$. The computations for the other cases in Table 1 run along the same lines. For $(x_{1}^{*},x_{2}^{*})\in\partial A_{1}^{*}$ and $\partial A_{3}^{*}$ the classification is simpler, since $\varphi(x_{1}^{*})-\varphi(x_{2}^{*})$ has constant sign.
@@
In the two examples shown in Figures 7 and 8, the trajectory goes near a folded node in the boundary of region $A_{3}$. The canard persists as a high frequency oscillation of small amplitude that alternates with the large amplitude relaxation oscillation of lower frequency in a mixed-mode oscillation. The small oscillations remain close to the synchrony plane while the large ones make alternate visits to the regions $A_{2}$ and $A_{4}$. The small oscillations take place close to a double fold point, that we proceed to discuss.
Figure 7. Mixed-mode oscillations arising from a canard on a solution of (2) near a folded node, parameters $b=0$, $c=-1.150079575$, $k=1$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(-1.5,2,\varphi(-1.5),\varphi(2)\right)$. (A) - Time course for $x_{1}(t)$ (blue) and $x_{2}(t)$ (orange). (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Projection of the trajectory (red) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the blue dot.
(D) - Projection of the trajectory (blue) on the $\left(x_{2},y_{1},y_{2}\right)$ space, initial condition on the blue dot, purple critical manifold $C_{0}$. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, conventions as in (B), white dot folded equilibrium. (F) - Time course for $y_{1}(t)$ (blue) and $y_{2}(t)$ (orange).
Figure 8. Mixed mode oscillations arising from a canard on a solution of (2) near a folded node, parameters $b=0$, $c=-1.1501075$, $k=0.5$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(-1.5,2,\varphi(-1.5),\varphi(2)\right)$. (A) - Time course for $x_{1}(t)$ (blue) and $x_{2}(t)$ (orange). (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Projection of the trajectory (blue) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the blue dot.
(D) - Projection of the trajectory (blue) on the $\left(x_{2},y_{1},y_{2}\right)$ space, initial condition on the blue dot, purple critical manifold $C_{0}$. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, conventions as in (B). (F) - Time course for $y_{1}(t)$ (blue) and $y_{2}(t)$ (orange).
@@ 5.2. Double fold points
In this section we deal with the dynamics around the double fold points $X=\left(x_{1}^{*},x_{2}^{*}\right)=\left(x^{*},\pm x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$. We start with the equations (6) and introduce a time rescaling of $\tau=t/\varphi^{\prime}(x_{1}^{*})\varphi^{\prime}(x_{2}^{*})$, that is singular at $x_{i}=x_{i}^{*}$, $i=1,2$. Then the equations (6) transform into
@@
Note that $F_{2}(x_{1},x_{2})=F_{1}(x_{2},x_{1})$.
Recall that we are using the notation $A_{i}^{*}$ for the sets $A_{i}^{*}=\left\{(x_{1},x_{2}):\ \left(x_{1},x_{2},\varphi(x_{1}),\varphi(x_{2})\right)\in A_{i}\right\}$, $i=1,\ldots,4$, we make the analogous convention for $S_{i}^{*}$, $i=1,\ldots,4$ and $R^{*}$ (see Figure 1). For $(x_{1},x_{2})\in A_{i}^{*}$ we have $\varphi^{\prime}(x_{j})<0$, $j=1,2$ hence the time rescaling preserves time orientation inside the $A_{i}^{*}$. Since $\varphi^{\prime}(x)>0$ for $-2/\sqrt{3}<x<2/\sqrt{3}$ then the time rescaling also preserves time orientation in $R^{*}$ and reverses it in the $S_{i}^{*}$, $i=1,\ldots,4$.
The main result of this section is the following:
@@
The parameters for the mixed-mode oscillations in Figures 7 and 8 satisfy condition (i). Note that condition (ii) only holds if $2/\sqrt{3}-(b+2k)\varphi(2/\sqrt{3})>0$ and that (iv) needs that $2/\sqrt{3}-(b+2k)\varphi(2/\sqrt{3})<0$.
@@
Let $F(x_{1},x_{2})=\left(F_{1}(x_{1},x_{2}),F_{2}(x_{1},x_{2})\right)$ and let $X=\left(x^{*},\pm x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$. Since $\varphi^{\prime}(x^{*})=0$, then $\dfrac{\partial F_{1}}{\partial x_{1}}(X)=\dfrac{\partial F_{2}}{\partial x_{2}}(X)=0$. Therefore, $\operatorname{Tr}DF(X)=0$ and hence, unless $\det DF(X)=0$, the point $X$ is either a saddle or a centre for the linearisation of the rescaled equations (13). A canard exists if there is a trajectory of (2) starting on the attracting part of the slow manifold $C_{\varepsilon}$ that crosses the fold line into the part of $C_{\varepsilon}$ that is not attracting. The idea of the proof is to show that under any of the conditions (i) and (ii) the point $X$ is a saddle and its stable manifold intersects the corresponding $A_{i}^{*}$, creating the possibility of canards near $X$, as in [25]. For conditions (iii) and (iv) the idea is to show that although the point $X$ is also a saddle, its stable manifold does not intersect the corresponding $A_{i}^{*}$, while its unstable manifold does. Hence canards are not possible around that point. We treat separately the two cases $X=\left(x^{*},x^{*}\right)$ and $X=\left(x^{*},-x^{*}\right)$.
@@
and hence, unless $c=x^{*}-b\varphi(x^{*})$, the point $X$ is a saddle. The eigenvalues of the matrix $DF(X)$ are $\lambda_{1}=\varphi^{\prime\prime}(x^{*})(x^{*}-b\varphi(x^{*})-c)$ and $\lambda_{2}=-\lambda_{1}$ with eigenspaces $V_{1}=\left\{\left(s,s\right)\ s\in{\mathbf{R}}\right\}$ and $V_{2}=\left\{\left(s,-s\right)\ s\in{\mathbf{R}}\right\}$, respectively. Thus locally one of the trajectories of (13) that is tangent to $V_{1}$ lies in $A_{1}^{*}$ or $A_{3}^{*}$, according to the case in question and all the trajectories of (13) that are tangent to $V_{2}$ lie outside $A_{1}^{*}$ and $A_{3}^{*}$.
@@
hence $\operatorname{Tr}DF(X)=0$, and $\det DF(X)=-{\varphi^{\prime\prime}(x^{*})}^{2}\left[(x^{*}-(b+2k)\varphi(x^{*}))^{2}-c^{2}\right]$. If $\det DF(X)>0$ then $DF(X)$ has a pair of purely imaginary eigenvalues. The point $X$ is a centre for the linearisation of the desingularised equations (13). This happens when $|c|>\left|(b+2k)\varphi(x^{*})-x^{*}\right|$, we do not include this situation in our analysis.
If $\det DF(X)<0$ then the point $X$ is a saddle, since the eigenvalues of $DF(X)$ are $\lambda_{\pm}=\pm\sqrt{-\det DF(X)}$. This happens if and only if
@@
Using the same arguments of the case $X=(x^{*},x^{*})$, it is possible to have canards if one branch of the stable manifold of $X$ lies in $A_{i}^{*}$, $i=2,4$. Canards are not possible if the condition fails. This is determined by the directions of the eigenspace associated to the negative eigenvalue $\lambda_{-}$, which is the set $V_{-}=\{(s(\alpha,\lambda_{-}):s\in{\mathbf{R}}\}$ where $\alpha=-\varphi^{\prime\prime}(x^{*})(x^{*}-(b+2k)\varphi(x^{*})-c)=\dfrac{\partial F_{1}}{\partial x_{2}}(X)$. There are no canards if $\alpha<0$, they are possible if $\alpha>0$.
@@
If $\beta>0$ then condition (16) is equivalent to $-\beta<\sigma c<\beta$ and this implies that $\alpha>0$. Since $\varphi(2/\sqrt{3})=16/3\sqrt{3}$ then $\beta>0$ if and only if $b+2k<3/8$.
If $\beta<0$ then condition (16) is equivalent to $\beta<\sigma c<-\beta$ and this implies that $\alpha<0$. Also $\beta<0$ if and only if $b+2k>3/8$.
We have established that $\alpha>0$ (implying canards are possblle) if and only if (16) holds and $\beta>0$. The necessary conditions are that $b+2k<3/8$ and $|c|<|\beta|=\beta$, as in assertion (ii). We have also established that $\alpha<0$ (implying canards are not posslble) if and only if (16) holds and $\beta<0$, and the necessary conditions are that $b+2k>3/8$ and $|c|<|\beta|=-\beta$, as in assertion (iv). ∎
@@
Figure 9. On the left, trajectories of (13) on the $\left(x_{1},x_{2}\right)$ plane, near the double fold point $X=\left(2/\sqrt{3},2/\sqrt{3}\right)\in\partial A_{1}^{*}$ for $\lambda_{1}<0$, in the rescaled time $\tau$. On the right, trajectories of (6) near the same point, in the original time $t$. The rescaling inverts the time orientation in regions $S_{j}$, $j=1,\ldots,4$ and preserves it in regions $R$ and $A_{j}$, $j=1,\ldots,4$. The dynamics of (6) is not well defined when $x_{i}=2/\sqrt{3}$, $i=1,2$.
@@
Let $X=\left(x^{*},x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$ be a double fold point. If $b=0$, $k>0$ and $\sigma c<2/\sqrt{3}$ then for $\sigma c$ close to $2/\sqrt{3}$ and small $\varepsilon>0$ there is a canard for (2) close to $X$.
@@
In Table 1 the conditions of the statement correspond to the first rows in the regions $\partial A_{1}^{*}$ and $\partial A_{3}^{*}$, where the point $\left(x_{1}^{*},x_{2}^{*}\right)$, with $x_{2}=x^{*}$ and $\sigma x_{1}^{*}>2/\sqrt{3}$ is either a node or a focus. The idea of the proof is to show that (a) as $c\to 2\sigma/\sqrt{3}$ these points accumulate on $X$, and that (b) for $\sigma c$ close to $2/\sqrt{3}$ the point is a node. Then it will follow, either by Lemma 2.3 of [25], or by the results of [12] and [13], that there are solutions of (2) starting close to $X$ in the attracting part of $C_{\varepsilon}$ that follow the repelling part of $C_{\varepsilon}$ for some time.

### v3

@@ 5.1. An open set of parameters where canards occur
which is not zero and less than 1/4 for an open set of parameters $(k,b,c)$, since $\varphi^{\prime\prime}(x_{2}^{*})=-6x_{2}^{*}\neq 0$ and $k\neq 0$. ∎
Figure 8. Illustration of Proposition 5.6. Canard transient on a solution of (2) near a folded node, parameters $b=0$, $c=-0.519935054$, $k=1$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(-1.5,2,{\varphi(-1.5),\varphi(2)}\right)$. (A) - Time course for $y_{1}(t)$ (blue) and $y_{2}(t)$ (orange). (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Projection of the trajectory (red) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the red dot, equilibria on the white dots. (D) - Projection of the trajectory (purple) on the $\left(x_{2},y_{1},y_{2}\right)$ space, initial condition on the blue dot, purple critical manifold $C_{0}$. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, conventions as in (B) – transient small oscillations near folded node. (F) - Time course for $x_{1}(t)$ (blue) and $x_{2}(t)$ (orange).
The next result is the goal of this subsection:
Proposition 5.6.
If $\left|x_{1}^{*}\right|>\dfrac{2}{\sqrt{3}}$ and $x_{2}^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$, there exists a non-empty open set of the parameters $(k,b,c)\in{\mathbf{R}}^{3}$ for which there is at least one trajectory of (2) that goes across $X^{*}=\left(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*})\right)$ generating a canard.
Proof.
First of all note that the time rescaling $\tau=t/\varphi^{\prime}(x_{2})$ we have used reverses time orientation when $\left(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*})\right)$ is in the attracting part of $C_{0}$ (note that here we have $\varphi^{\prime}(x_{2}^{*})<0$). Therefore when $(x_{1}^{*},x_{2}^{*})$ is either a saddle or an unstable node for (7) there is at least one trajectory of (2) that goes across $\left(x_{1}^{*},x_{2}^{*},\varphi(x_{1}^{*}),\varphi(x_{2}^{*})\right)$ generating a canard. This is a consequence of Lemma 5.4, where it is stated that canards correspond to, at least, one negative eigenvalue of $DH$ associated to (7) at a fold point. ∎
Examples are shown in Figures 8, 9 and 10. In the first of these examples, before the trajectory approaches a synchronous periodic orbit it makes some transient small oscillations around the fold point (“canard behaviour-type”) that are visible in Figure 8 (E). In this example there is a folded node, see Table 1 below.
Note that from equation (10) it follows that for any given $k\neq 0$ and $b\in{\mathbf{R}}$ there is a value of the parameter $c\in{\mathbf{R}}$ for which $\det DH(x_{1}^{*},x_{2}^{*})=0$. This corresponds to saddle-node bifurcations. According to [23] this gives rise to mixed-mode oscillations like those shown in Figures 9 and 10, i.e., trajectories that combine small oscillations and large oscillations of relaxation type, both recurring in an alternating manner.
@@ 5.2. The case $b=0$
As an illustration of Lemmas 5.2 and 5.5 we describe their contents in a particular case. More information on the equilibria of the equation (7) may be obtained if we make the simplifying assumption $b=0$. Indeed the expression of (1) in the particular case $b=0$ has been used in the analysis of two FHN coupled in the fast equations by Pedersen et al. [29], in experimental results with an electrical circuit by Kulminskiy et al. [26] and with double coupling fast–to–fast and slow–to–slow by Krupa et al. 2014 [23]. Thus this particular case is interesting for comparing the outcomes of different types of couplings. We obtain sufficient conditions on the parameters $c$ and $k\neq 0$ for the existence of a folded saddle. Since by Lemma 5.5 this holds in an open set of parameters, then the folded saddles will persist for small $b\neq 0$.
@@
$A_{1}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2})\in{\mathbf{R}}^{4}:\ x_{1}>2/\sqrt{3},\quad x_{2}>2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}$
$A_{2}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2})\in{\mathbf{R}}^{4}:\ x_{1}<-2/\sqrt{3},\quad x_{2}>2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}$
$A_{3}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2})\in{\mathbf{R}}^{4}:\ x_{1}<-2/\sqrt{3},\quad x_{2}<-2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}$
$A_{4}$ $=$ $\left\{(x_{1},x_{2},y_{1},y_{2})\in{\mathbf{R}}^{4}:\ x_{1}>2/\sqrt{3},\quad x_{2}<-2/\sqrt{3},\quad y_{i}=\varphi(x_{i}),\quad i=1,2\right\}.$
@@
In Table 1, we present sufficient conditions on $c$ for the existence and classification of a folded equilibrium of (7) with $b=0$ on the components with $x_{2}^{*}=\pm 2/\sqrt{3}$ of $\partial A_{i}^{*}$, $i=1,\ldots,4$, the boundary of the attracting region of the critical manifold. The analysis for folded equilibria with $x_{1}^{*}=\pm 2/\sqrt{3}$ follows by symmetry. We present the computations for the cases where the folded equilibrium $\left(x_{1}^{*},x_{2}^{*}\right)$ lies on the boundaries of $A_{1}^{*}$ or $A_{2}^{*}$. The computations for the other cases in Table 1 run along the same lines. For $(x_{1}^{*},x_{2}^{*})\in\partial A_{1}^{*}$ or $\partial A_{3}^{*}$ the classification is simpler, since $\varphi(x_{1}^{*})-\varphi(x_{2}^{*})$ has constant sign.
@@
In the two examples shown in Figures 9 and 10, the trajectory goes near a folded node in the boundary of region $A_{3}$. The canard persists as a high frequency oscillation of small amplitude that alternates with the large amplitude relaxation oscillation of lower frequency in a mixed-mode oscillation. The small oscillations remain close to the synchrony plane while the large ones make alternate visits to the regions $A_{2}$ and $A_{4}$. The small oscillations take place close to a double fold point, that we proceed to discuss.
Figure 9. Illustration of Theorem 5.8. Mixed-mode oscillations arising from a canard on a solution of (2) near a folded node, parameters $b=0$, $c=-1.150079575$, $k=1$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(-1.5,2,\varphi(-1.5),\varphi(2)\right)$. These numerics illustrate the combination of large and small oscillations in a persistent and regular way. (A) - Time course for $x_{1}(t)$ (blue) and $x_{2}(t)$ (orange). (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Projection of the trajectory (red) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the blue dot. (D) - Projection of the trajectory (blue) on the $\left(x_{2},y_{1},y_{2}\right)$ space, initial condition on the blue dot, purple critical manifold $C_{0}$. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, conventions as in (B), white dot folded equilibrium. (F) - Time course for $y_{1}(t)$ (blue) and $y_{2}(t)$ (orange).
Figure 10. Illustration of Theorem 5.8. Mixed mode oscillations arising from a canard on a solution of (2) near a folded node, parameters $b=0$, $c=-1.1501075$, $k=0.5$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(-1.5,2,\varphi(-1.5),\varphi(2)\right)$. (A) - Time course for $x_{1}(t)$ (blue) and $x_{2}(t)$ (orange). (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Projection of the trajectory (blue) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the blue dot. (D) - Projection of the trajectory (blue) on the $\left(x_{2},y_{1},y_{2}\right)$ space, initial condition on the blue dot, purple critical manifold $C_{0}$. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, conventions as in (B). (F) - Time course for $y_{1}(t)$ (blue) and $y_{2}(t)$ (orange).
@@ 5.3. Double fold points
lixo
The treatment of folded equilibria of Subsection 5.1 above cannot be applied around the double fold points where two fold lines cross. This is because at these points both $\varphi^{\prime}(x_{1}^{*})=0$ and $\varphi^{\prime}(x_{2}^{*})=0$ and hence the first equation in (7) is not well defined, so these points have to be analysed separately. In this section we deal with the dynamics around the double fold points $X=\left(x_{1}^{*},x_{2}^{*}\right)=\left(x^{*},\pm x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$.
We start with the equations (6) and introduce a time rescaling of $\tau=t/\varphi^{\prime}(x_{1}^{*})\varphi^{\prime}(x_{2}^{*})$, that is singular at $x_{i}=x_{i}^{*}$, $i=1,2$. Then the equations (6) transform into
@@
Note that $F_{2}(x_{1},x_{2})=F_{1}(x_{2},x_{1})$. In what follows, let $F(x_{1},x_{2})=\left(F_{1}(x_{1},x_{2}),F_{2}(x_{1},x_{2})\right)$, be defined in $(x_{1},x_{2})\in{\mathbf{R}}^{2}\backslash\{(x_{1},x_{2}):\varphi^{\prime}(x_{1})\varphi^{\prime}(x_{2})=0\}$.
Recall that we are using the notation $A_{i}^{*}$ for the sets
$A_{i}^{*}=\left\{(x_{1},x_{2})\in{\mathbf{R}}^{2}:\ \left(x_{1},x_{2},\varphi(x_{1}),\varphi(x_{2})\right)\in A_{i}\right\},\quad i=1,\ldots,4,$
we make the analogous convention for $S_{i}^{*}$, $i=1,\ldots,4$ and $R^{*}$ (see Figure 2). For $(x_{1},x_{2})\in A_{i}^{*}$ we have $\varphi^{\prime}(x_{j})<0$, $j=1,2$ hence the time rescaling preserves time orientation inside the $A_{i}^{*}$. Since $\varphi^{\prime}(x)>0$ for $-2/\sqrt{3}<x<2/\sqrt{3}$ then the time rescaling also preserves time orientation in $R^{*}$ and reverses it in the $S_{i}^{*}$, $i=1,\ldots,4$. The main result of this section is the following:
@@
The parameters for the mixed-mode oscillations in Figures 9 and 10 satisfy condition (i). Note that condition (ii) only holds if $2/\sqrt{3}-(b+2k)\varphi(2/\sqrt{3})>0$ and that (iv) needs that $2/\sqrt{3}-(b+2k)\varphi(2/\sqrt{3})<0$.
A canard exists if there is a trajectory of (2) starting on the attracting part of the slow manifold $C_{\varepsilon}$ that crosses the fold line into the part of $C_{\varepsilon}$ that is not attracting. The idea of the proof of Theorem 5.7 is to show that under any of the conditions (i) and (ii) the point $X$ is a saddle for $F$ and its stable manifold intersects the corresponding $A_{i}^{*}$, creating the possibility of canards near $X$, as in [34, Lemma 3.1]. For conditions (iii) and (iv) the idea is to show that although the point $X$ is also a saddle, its stable manifold does not intersect the corresponding $A_{i}^{*}$, while its unstable manifold does. Hence canards are not possible around that point.
@@
Let $X=\left(x^{*},\pm x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$. Since $\varphi^{\prime}(x^{*})=0$, then $\dfrac{\partial F_{1}}{\partial x_{1}}(X)=\dfrac{\partial F_{2}}{\partial x_{2}}(X)=0$. Therefore, $\operatorname{tr}DF(X)=0$ and hence, unless $\det DF(X)=0$, the point $X$ is either a saddle or a centre for the linearisation of the rescaled equations (13). We treat separately the two cases $X=\left(x^{*},x^{*}\right)$ and $X=\left(x^{*},-x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$.
First case
@@
and hence, unless $c=x^{*}-b\varphi(x^{*})$, the point $X$ is a saddle. The eigenvalues of the matrix $DF(X)$ are $\lambda_{1}=\varphi^{\prime\prime}(x^{*})(x^{*}-b\varphi(x^{*})-c)$ and $\lambda_{2}=-\lambda_{1}$ with eigenspaces
$V_{1}=\left\{\left(s,s\right)\ s\in{\mathbf{R}}\right\}\quad\text{and}\quad V_{2}=\left\{\left(s,-s\right)\ s\in{\mathbf{R}}\right\}$
, respectively. Thus locally one of the trajectories of (13) that is tangent to $V_{1}$ lies in $A_{1}^{*}$ or $A_{3}^{*}$, according to the case in question and all the trajectories of (13) that are tangent to $V_{2}$ lie outside $A_{1}^{*}$ and $A_{3}^{*}$.
@@
hence $\operatorname{tr}DF(X)=0$, and $\det DF(X)=-{\varphi^{\prime\prime}(x^{*})}^{2}\left[(x^{*}-(b+2k)\varphi(x^{*}))^{2}-c^{2}\right]$. If $\det DF(X)>0$, then $DF(X)$ has a pair of purely imaginary eigenvalues. The point $X$ is a centre for the linearisation of the desingularised equations (13). This happens when $|c|>\left|(b+2k)\varphi(x^{*})-x^{*}\right|$, we do not include this situation in our analysis.
If $\det DF(X)<0$, then the point $X$ is a saddle, since the eigenvalues of $DF(X)$ are $\lambda_{\pm}=\pm\sqrt{-\det DF(X)}$. This happens if and only if
@@
Using the same arguments of the case $X=(x^{*},x^{*})$, it is possible to have canards if one branch of the stable manifold of $X$ lies in $A_{i}^{*}$, $i=2,4$. Canards are not possible if the condition fails. This is determined by the directions of the eigenspace associated to the negative eigenvalue $\lambda_{-}$, which is the set $V_{-}=\{(s(\alpha,\lambda_{-}):s\in{\mathbf{R}}\}$ where
$\alpha=-\varphi^{\prime\prime}(x^{*})(x^{*}-(b+2k)\varphi(x^{*})-c)=\dfrac{\partial F_{1}}{\partial x_{2}}(X).$
There are no canards if $\alpha<0$, they are possible if $\alpha>0$.
@@
If $\beta>0$, then condition (16) is equivalent to $-\beta<\sigma c<\beta$ and this implies that $\alpha>0$. Since $\varphi(2/\sqrt{3})=16/3\sqrt{3}$ then $\beta>0$ if and only if $b+2k<3/8$.
If $\beta<0$, then condition (16) is equivalent to $\beta<\sigma c<-\beta$ and this implies that $\alpha<0$. Also $\beta<0$ if and only if $b+2k>3/8$.
We have established that $\alpha>0$ (implying canards are possible) if and only if (16) holds and $\beta>0$. The necessary conditions are that $b+2k<3/8$ and $|c|<|\beta|=\beta$, as in assertion (ii). We have also established that $\alpha<0$ (implying canards are not possible) if and only if (16) holds and $\beta<0$, and the necessary conditions are that $b+2k>3/8$ and $|c|<|\beta|=-\beta$, as in assertion (iv). ∎
@@
Figure 11. On the left, trajectories of (13) on the $\left(x_{1},x_{2}\right)$ plane, near the double fold point $X=\left(2/\sqrt{3},2/\sqrt{3}\right)\in\partial A_{1}^{*}$ for $\lambda_{1}<0$, in the rescaled time $\tau$. On the right, trajectories of (6) near the same point, in the original time $t$. The rescaling inverts the time orientation in regions $S_{j}$ (green), $j=1,\ldots,4$ and preserves it in regions $R$ (blue) and $A_{j}$ (yellow), $j=1,\ldots,4$. The dynamics of (6) is not well defined when $x_{i}=2/\sqrt{3}$, $i=1,2$.
@@
Let $X=\left(x^{*},x^{*}\right)$, $x^{*}=2\sigma/\sqrt{3}$, $\sigma=\pm 1$ be a double fold point. If $b=0$, $k>0$ and $\sigma c<2/\sqrt{3}$, then for $\sigma c$ close to $2/\sqrt{3}$ and small $\varepsilon>0$ there is a canard for (2) close to $X$.
@@
In Table 1 the conditions of the statement correspond to the first rows in the regions $\partial A_{1}^{*}$ and $\partial A_{3}^{*}$, where the point $\left(x_{1}^{*},x_{2}^{*}\right)$, with $x_{2}=x^{*}$ and $\sigma x_{1}^{*}>2/\sqrt{3}$ is either a node or a focus. The idea of the proof is to show that:
  1. (a)
    as $c\to 2\sigma/\sqrt{3}$ these points accumulate on $X$, and
  2. (b)
    for $\sigma c$ close to $2/\sqrt{3}$ the point is a node.
Then it will follow, either by Lemma 2.3 of [34], or by the results of [18] and [19], that there are solutions of (2) starting close to $X$ in the attracting part of $C_{\varepsilon}$ that follow the repelling part of $C_{\varepsilon}$ for some time.